OpenStudy (anonymous):

What is the sum of: 12/z-5 + 12/z+4

OpenStudy (anonymous):

Hint: find a common denominator

OpenStudy (anonymous):

umm.. 20?

OpenStudy (anonymous):

Wait, so it's not 12/(z-5)? It's 12/z -5?

OpenStudy (anonymous):

\[\frac{ 12 }{ z-5 } + \frac{ 2 }{ z+4 }\]

OpenStudy (anonymous):

Common denominator is not 20. You've got a somewhat idea in that you multiply but you're not multiplying the whole denominator.

OpenStudy (anonymous):

I don't understand..... sorry, can you, like, re-word it, i'm kinda slow..

OpenStudy (anonymous):

Just multiply the denominators to find common denominator

OpenStudy (anonymous):

so i have to multiply z-5 and z+4?

OpenStudy (anonymous):

Common denominator would be (z-5)(z+4)

OpenStudy (anonymous):

then the numerator would be 14?

OpenStudy (anonymous):

No if you multiply something to bottom, you multiply the same thing to the top or else you change the entire thing

OpenStudy (anonymous):

So what would u multiply to 12?

OpenStudy (anonymous):

oh, so is the answer, \[\frac{ 2(7z+19) }{(z-5)(z+4) }\]?

OpenStudy (anonymous):

No (z+4) because ur multiplying the bottom by (z+4)

OpenStudy (anonymous):

so, wait, the denominator is (z-5)(z+4)?

OpenStudy (anonymous):

Yes, that's ur common denominator. Let me draw it out for u

OpenStudy (anonymous):

would it be better if i give you the multiple choice answers?

OpenStudy (anonymous):

12(z+4) +2(z-5) /(z+4)(z-5)

OpenStudy (anonymous):

nothing your saying is one of the answers

OpenStudy (anonymous):

14z+38 /common den.

OpenStudy (anonymous):

14z + 38 / (z+4)(z-5)

OpenStudy (anonymous):

that's not one of the answers..

OpenStudy (anonymous):

24z-12 / common deno

OpenStudy (anonymous):

12/(z-5) + 12/(z+4) 12 * (z+4) = 12z + 48 12 * (z-5) = 12z - 60 So 12z + 48 + 12z - 60 = 24z - 12.

OpenStudy (anonymous):

24z - 12/(z-5)(z+4) OR 12(2z -1)/(z-5)(z+4). Either one should be the answer.

OpenStudy (anonymous):